The molecular liquid dichloroethane (C2H4C) has a density of 1.25 gmL, while the molecular liquid dibromoethane (C2H,Br2) has a density of 2.17 gmL (a) When 69.1 mL of C,H4C2 and 38.9 mL of C,H4Br2 are mixed, a clear liquid solution forms. Is this a chemical or a physical change? Is the solution a pure substance, a homogeneous mixture or a heterogeneous mixture? (b) Is the density of this solution a chemical or a physical property? (c) Assuming that the volumes are additive upon mixing, what is the density of the solution? g mL
Answer #1

(a) Both of dichloro ethane and dibromomethane are colorless liquids, on mixing a clear liquid solution is formed

Since the clear liquid form which means it is colorless, hence there is no color change in the experiment, so this is a PHYSICAL CHANGE

It is not a pure substance, since there are two liquids mixed but they have formed a clear solution and mixed with each other, so it is a HOMOGENEOUS MIXTURE

b) Density of the compound is a physical property, since it only depends on the mass and volume of the substance. We don't require any reaction or chemical change to find density

Hence density is a PHYSICAL property

c)

Using the formula

Total volume = V1 + V2 = 69.1 + 38.9 = 108 mL

d1v1 + d2v2 = d * (v1+v2)

69.1 * 1.25 + 38.9 * 2.17 = 108 * d

d = 1.58 g/mL